Translate the following into an equation, and then solve the equation.
Five times a number decreased by three is 17. Determine the number.

Answer by addingup(3677)

Five times a number: 5x
Decreased by 3: -3
5*x-3 = 17
5x = 20
5x/5 = 20/5
x = 4